OpenStudy (anonymous):

Please help me! I don't get this ! Can someone explain how to do it? Thanks! Write the equation of the line that is parallel to the line y = −3x + 12 and passes through the point (−1, 6). y = 1/3x + 7 y = −3x + 3 y = 1/3x + 3 y = −3x + 7

OpenStudy (angela210793):

the line must be parallel with y = −3x + 12 .Two lines are parallel when their slopes are the same so the slope of the line we are looking for must be -3. when we got the slope and a point we use this formula: y-y1=slope(x-x1) so y-6=-3(x+1) y=-3x+3

OpenStudy (angela210793):

Get it Mimi?

OpenStudy (anonymous):

I think... does the same thing go for perpendicular lines?

OpenStudy (angela210793):

no the slope of the line when it is perpendicular is -1/slope of the line given then it's the same :) :)

OpenStudy (anonymous):

Okay, thanks a bunch! :)

OpenStudy (angela210793):

yw :)